OpenStudy (anonymous):

Differentiate the function f(z)=(1-1/z)^5

OpenStudy (perl):

first rewrite the expression f(z) = ( 1 - z^-1 ) ^5 f ' (z) = 5 ( 1 - z^ -1 ) ^ 4 * ( -(-1)*z^-2))

OpenStudy (anonymous):

Simplify and apply quotient rule

OpenStudy (perl):

you can use chain rule

OpenStudy (anonymous):

\[\frac{ d }{ dx } u^n = n u ^{n-1}u'\]

OpenStudy (anonymous):

Where in your case u = the function inside ( ) and u' is the derivative of that function

OpenStudy (perl):

f ' (z) = 5 ( 1 - z^ -1 ) ^ 4 * ( -(-1)*z^-2)) f ' (z) = 5 ( 1 - z^ -1 ) ^ 4 * ( z^-2)) f ' (z) = 5 ( 1 - z^ -1 ) ^ 4 / ( z^2) f ' (z) = 5 ( 1 - 1/z ) ^4 / z^2